# VideoLLM Knows When to Speak: Enhancing Time-Sensitive Video Comprehension with Video-Text Duet Interaction Format

### v1

@@ 6.2 Dense Video Captioning
[…]
MMDuet (Ours)
$t=0.6$usym2714 2.46/2.33 13.7/4.0 1.80
$t=0.5$usym2714 2.77/2.61 18.4/5.3 2.23
$t=0.4$usym2714 3.00/2.81 23.0/6.6 2.59
$t=0.3$usym2714 3.13/2.93 27.0/7.6 2.73
Table 4: Results on the test set of Shot2Story-MAGQA-39k with the rm. ass. turns method used. For the “time per example” column, the time used by “LLaVA-OV-VT” is set to 1, and the times for other rows are set as multiples of the time used by “LLaVA-OV-VT”.
6.3 Multi-Answer Grounded Video QA
To align closely with the widely-used streaming video comprehension scenario, we propose MAGQA that requires a model to generate answers at multiple necessary positions of a video. Different from conventional Video QA in which one question corresponds to only one answer, In MAGQA, a question corresponds to multiple turns of answers, and these turns are derived from different video segments. Therefore, this task requires the response to be accurate and in time.
We introduce an “in-span score” metric, which takes both the text content and reply time of the predicted answer into account to evaluate the model’s performance on this task. Suppose the model prediction has $P$ answers, each answer has a prediction time $time_{p}$ and prediction text $pred_{p}$, $p=1,2,\dots,P$. The ground truth has $Q$ answers, each answer has a ground truth start time $start_{q}$, a ground truth end time $end_{q}$, and a ground truth text $gold_{q}$, $q=1,2,\dots,Q$. First, we use an LLM to calculate a relevance score from 1 to 5 between each answer in prediction $pred_{p}$ and ground truth $gold_{q}$: $S=\{s_{p,q}\}\in\mathcal{R}^{P\times Q}$. For each ground truth answer $q$, we select the predicted answers with predicted time in ground truth time span: $\mathcal{P}_{q}=\{p\mid time_{p}\in[start_{q},end_{q}]\}$, and use the average score between the ground truth answer and the selected predicted answers as the score for this ground truth answer: $score_{q}=\frac{1}{|\mathcal{P}_{q}|}\sum_{p\in\mathcal{P}_{q}}s_{p,q}\quad\text{if }|\mathcal{P}_{q}|>0$. If $|\mathcal{P}_{q}|=0$ (no predicted answer falls in this ground truth span), $score_{q}$ is set to $1$. Finally, we calculate the average score of all ground truth answers as the final in-span score of this example: $in\_span\_score=\frac{1}{|Q|}\sum_{q=1}^{|Q|}score_{q}$. GPT-4o-2024-08-06 [20] is used as the scorer LLM for in-span score. To prevent reproducibility issues due to potential changes of OpenAI API, we also report the in-span score obtained using LLaMA 3.1 70B Instruct [2].
As MAGQA requires the answers to be both informative and related to the question, we set the need_response as: if the sum of informative score and relevance score of a frame is larger than a threshold $t$, then the model needs to generate a response right after this frame. We also use the “rm. prev. resp.” method in dense video captioning task introduced in Sec. 6.2. As baseline models are not capable of generating responses at specific positions in the video, we employ an output format the same as dense video captioning, i.e., output the start time, end time, and predicted text for each turn after watching the entire video in both training and testing, and use the average of the start and end time as the response time. For the cases that baseline models directly reply with an answer without corresponding time, we pair this predicted answer with all ground truth answers. Note that this is a simplified requirement than that of MMDuet, as the MAGQA task simulates streaming video comprehension application scenario, which requires the model to respond as soon as the video plays to segments relevant to the question, which ensures that users can see the responses timely, rather than waiting until the entire video concludes before generating replies.
Results are shown Tab. 4. We provide results for different $t$ as it represents a trade-off between inference time and performance: as $t$ decreases from $0.6$ to $0.3$, the performance of MMDuet’s real-time replies continuously rise and even outperform baselines with a simplified setting of providing non-real-time replies after watching the entire video. However, this is achieved at a cost of generating lots of duplicate replies with around 2.5$\times$ inference time.
Model YouCook2
MMDuet 2.9/8.8/21.7
w/o rand. resp. pos. 2.1/7.3/19.0
w/o multi informative 2.9/8.0/16.5
Table 5: Ablation study on training methods.
6.4 Ablation Studies
We conduct ablation studies on YouCook2 dense video captioning to assess two empirical yet important findings for effectively training the informative head in data construction: randomly inserting the response at a position from 50% to 75% of the corresponding video segment (rand. resp. pos.), and setting informative head’s label to TRUE for all frames between 50% of the segment and the response time (multi informative). When “rand. resp. pos.” is disabled, the response is always inserted at the end of the corresponding segment. When “multi informative” is disabled, only the informative label of the frame right before the response is set as TRUE. As illustrated in Tab. 5, disabling either method negatively impact MMDuet’s performance, which shows the importance of carefully handling the response time and informative labels.
7 Limitations and Future Works
As a preliminary and pioneering effort of video-text duet interaction format, we acknowledge that there is much room for improvement which should be addressed in future research: (1) Some hyperparameters (e.g., the need_response criterion) are required during inference. Though the parameter sensitivity test shows that this criterion is quite robust across different thresholds, there still exists unknowns for applying MMDuet on new tasks. (2) Information from subsequent frames is not incorporated when generating in-time responses for the current frame, especially for the live-streaming video that indeed has unpredictable future frames. It can be crucial in some scenarios, such as determining the start of an action. (3) Slow inference speed. A better inference process is needed to improve the parallelism of video inputs and avoid generating duplicate responses. (4) Real-time response datasets with longer live-streaming videos are required to be collected to better fit the real-world application scenarios.
8 Conclusion
In this paper, we formalize the video-text duet interaction format, which addresses the shortcomings of existing interaction formats in terms of time-sensitive task performance, more application scenarios and in-time information retrieval for long videos. We collect MMDuetIT, a dataset for training models to follow the video-text duet interaction format. Based on MMDuetIT we train MMDuet, a model with significant improvements on various time-sensitive tasks, and is able to generate responses in real time. We believe such improvements can be a substantial step towards building powerful and useful video comprehension systems.
Supplementary Material
Appendix A Details of Training MMDuet
A.1 Training Hyperparameters.
LLaVA-OneVision uses SigLIP-Large [39] as the vision encoder, and converts an image with $384\times 384$ into $24\times 24=576$ tokens. In the official settings of LLaVA-OneVision [15], when encoding videos, the visual tokens corresponding to each frame are spatially downsampled to $12\times 12=144$ tokens using a pooling operation with a size of 2. However, this number of tokens is also too large when training and inference with long videos. To address this, we further modified the pooling size to 4, resulting in $7\times 7=49$ tokens per frame.
[…]
@@ Appendix A Details of Training MMDuet
The projector, the relevance head, the informative head and LoRA [9] weights of the LLM (add to all attention proj. layers and FFN layers) are trained, while other parameters of the model are frozen. More training hyperparameters are listed in Tab. 6.
A.2 Inference Settings
[…]
@@ A.2 Inference Settings
Appendix B Example Inputs for Each Task in MMDuetIT
Example inputs for each task for training and inference are listed in Tab. 7. The dense video captioning user input is selected from one of the following sentences:
@@ Appendix B Example Inputs for Each Task in MMDuetIT
Dense Video Captioning <im_start>system A multimodal AI assistant is helping users with some activities. Below is their conversation, interleaved with the list of video frames received by the assistant. <im_end> <im_start>user (A Dense Video Captioning Query)<im_end> <im_start>stream <frame><frame><frame> …<im_end> <im_start>assistant A person pulls a knife from a black bag.<im_end> <im_start>stream <frame><frame><frame> …<im_end> <im_start>assistant A man in a hat and red clothes speaks with a dagger, and a tree behind him.<im_end> <im_start>stream <frame><frame><frame> …<im_end> …(More stream and assistant turns)
MAGQA <im_start>system A multimodal AI assistant is helping users with some activities. Below is their conversation, interleaved with the list of video frames received by the assistant. <im_end> <im_start>stream <frame><frame><frame> …<im_end> <im_start>user What happens during the basketball game?<im_end> <im_start>stream <frame><frame><frame> …<im_end> <im_start>assistant Several players in white jerseys are celebrating by high-fiving each other.<im_end> <im_start>stream <frame><frame><frame> …<im_end> <im_start>assistant A player in a white jersey makes a successful shot.<im_end> <im_start>stream <frame><frame><frame> …<im_end> …(More stream and assistant turns)
Temporal Video Grounding <im_start>system A multimodal AI assistant is helping users with some activities. Below is their conversation, interleaved with the list of video frames received by the assistant. <im_end> <im_start>user (A Temporal Video Grounding Query)<im_end> <im_start>stream <frame><frame><frame> …<im_end>
@@
Appendix C Qualitative Study
We list some examples of dense video captioning on videos with several minutes in length and contains many actions in Figs. 6, 7 and 8, and examples of multi-answer grounding video question answering (MAGQA) in Figs. 9, 10 and 11. For LLaVA-OV-TC and LLaVA-OV-VT, we directly list their generated outputs. For MMDuet, we list the numerical order (in round brackets), time (in square brackets) and content (in the second line) for each turn. If a line contains multiple numerical orders and times, this indicates that these turns have the same content, which is shown in the following line. To help readers to identify the position of these turns within the video, we also annotate the numerical order of the turns at the corresponding timestamps in the video stream.

### v2

@@ B.2 Training Hyperparameters
The projector, the relevance head, the informative head and LoRA hu2022lora weights of the LLM (add to all attention proj. layers and FFN layers) are trained, while other parameters of the model are frozen. More training hyperparameters are listed in Table˜5.
B.3 Pseudo Code of the Inference Process
We provide a python-style pseudo code of the inference process in LABEL:alg:inference.
1# Input:
2# system_prompt
3# video: list of frames
4# fps: frames per second to sample from video
5# user_turns: list of (time, text) sorted by time
6# Output:
7# model_turns: generated list of (time, text)
8
9model_turns = []
10v_inf_list, v_rel_list = [], []
11kv_cache = model(system_prompt)
12time = 0
13for frame in video:
14 if len(user_turns) and time>=user_turns[0].time:
15 kv_cache = model(kv_cache, user_turns[0].text)
16 user_turns = user_turns[1:]
17 kv_cache, v_inf, v_rel = model(kv_cache, frame)
18 v_inf_list.append(v_inf) # informative score
19 v_rel_list.append(v_rel) # relevance score
20 if need_response(v_inf_list, v_rel_list):
21 kv_cache, response = model.generate(kv_cache)
22 model_turns.append((time, response))
23 time += 1 / fps
Listing 1: Inference Process of MMDuet
B.4 Inference Settings
Videos from different sources are also sampled with different fps during inference. Specifically, we set the maximum number of frames sampled from each video to 400, and fps to 2 for videos from Shot2Story Han2023Shot2Story20KAN and Charades-STA Gao2017TALLTA, 1 for videos from QVHighlights Lei2021QVHighlightsDM, and 0.5 for videos from YouCook2 Zhou2017TowardsAL. For a few videos in YouCook2 that are even longer than $400\text{(frames)}\div 0.5\text{(fps)}=800$ seconds, we uniformly sample $400$ frames from this video to ensure that information from the latter part of the video is not truncated. This inference setting is consistent across MMDuet, LLaVA-OV-TC, and LLaVA-OV-VT.
@@ B.4 Inference Settings
B.5 Details of the In-Span Score
Suppose the model prediction has $P$ answers, each answer has a prediction time $time_{p}$ and prediction text $pred_{p}$, $p=1,2,\dots,P$. The ground truth has $Q$ answers, each answer has a ground truth start time $start_{q}$, a ground truth end time $end_{q}$, and a ground truth text $gold_{q}$, $q=1,2,\dots,Q$. First, we use an LLM to calculate a relevance score from 1 to 5 between each answer in prediction $pred_{p}$ and ground truth $gold_{q}$: $S=\{s_{p,q}\}\in\mathcal{R}^{P\times Q}$. For each ground truth answer $q$, we select the predicted answers with predicted time in ground truth time span: $\mathcal{P}_{q}=\{p\mid time_{p}\in[start_{q},end_{q}]\}$, and use the average score between the ground truth answer and the selected predicted answers as the score for this ground truth answer: $score_{q}=\frac{1}{|\mathcal{P}_{q}|}\sum_{p\in\mathcal{P}_{q}}s_{p,q}\quad\text{if }|\mathcal{P}_{q}|>0$. If $|\mathcal{P}_{q}|=0$ (no predicted answer falls in this ground truth span), $score_{q}$ is set to $1$. Finally, we calculate the average score of all ground truth answers as the final in-span score of this example: $in\_span\_score=\frac{1}{|Q|}\sum_{q=1}^{|Q|}score_{q}$.
Appendix C More Experimental Results
C.1 Hyperparameter Sensitivity
We list the experiments using different window size $w$ for temporal grounding in Fig.˜4 and threshold $s$ for dense captioning in Fig.˜5.
Figure 5: Performance on dense video captioning with different $s$.
C.2 Details of the Proactive Output Experiment
More results and baselines are listed in Table˜6. For results of models without streaming abilities (Proprietary MLLMs & Open-Sourced VideoLLMs), we follow the evaluation method of Lin2024StreamingBenchAT and Qian2025DispiderEV: We gradually extend the input video one second at a time and ask the model with the question “Is it the right time to output?”. If the model responds with “Yes.”, this moment is recorded as the predicted output timestamp. For MMDuet, we use the time of the first reply after the user question is input as the predicted output timestamp. For examples that MMDuet does not provide any reply at all, we consider them as failing cases and the difference between ground truth output time and predicted output time is recorded as $\infty$.
Model Acc Model Acc
Proprietary MLLMs
Gemini 1.5 pro 45.10 GPT-4o 56.86
Claude 3.5 Sonnet 64.71
Open-Sourced VideoLLMs
LLaVA-OV 29.55 Qwen2-VL 22.73
MiniCPM-V 2.6 22.22 LLaVA-NeXT-Video 18.18
InternVL2 40.91 LongVA 15.91
Streaming MLLMs
Flash-VStream 1.96 VideoLLM-Online 3.92
Dispider 25.34
MMDuet $t=0.3$ 29.44 MMDuet $t=0.4$ 31.85
MMDuet $t=0.5$ 26.61 MMDuet $t=0.6$ 18.95
Table 6: Performance of more baselines and MMDuet on the Proactive Output task of StreamingBench with different $t$.
Appendix D Example Inputs for Each Task in MMDuetIT
Example inputs for each task for training and inference are listed in Table˜7. The dense video captioning user input is selected from one of the following sentences:
@@ Appendix D Example Inputs for Each Task in MMDuetIT
Dense
Video
Captioning
<im_start>system
A multimodal AI assistant is helping users with some activities. Below is their conversation, interleaved with the list of video frames received by the assistant. <im_end>
<im_start>user
(A Dense Video Captioning Query)<im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
<im_start>assistant
A person pulls a knife from a black bag.<im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
<im_start>assistant
A man in a hat and red clothes speaks with a dagger, and a tree behind him.<im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
…(More stream and assistant turns)
MAGQA
<im_start>system
A multimodal AI assistant is helping users with some activities. Below is their conversation, interleaved with the list of video frames received by the assistant. <im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
<im_start>user
What happens during the basketball game?<im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
<im_start>assistant
Several players in white jerseys are celebrating by high-fiving each other.<im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
<im_start>assistant
A player in a white jersey makes a successful shot.<im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
…(More stream and assistant turns)
Temporal
Video
Grounding
<im_start>system
A multimodal AI assistant is helping users with some activities. Below is their conversation, interleaved with the list of video frames received by the assistant. <im_end>
<im_start>user
(A Temporal Video Grounding Query)<im_end>
<im_start>stream
<frame><frame><frame> …<im_end>
@@
Appendix E Qualitative Study
We list some examples of dense video captioning on videos with several minutes in length and contains many actions in Figs.˜6, 7 and 8, and examples of multi-answer grounding video question answering (MAGQA) in Figs.˜9, 10 and 11. For LLaVA-OV-TC and LLaVA-OV-VT, we directly list their generated outputs. For MMDuet, we list the numerical order (in round brackets), time (in square brackets) and content (in the second line) for each turn. If a line contains multiple numerical orders and times, this indicates that these turns have the same content, which is shown in the following line. To help readers to identify the position of these turns within the video, we also annotate the numerical order of the turns at the corresponding timestamps in the video stream.